If we run a one-sample t-test and get a t-stat=6.74, with n=7, the p-value under one-sided conditions is:
Answer #1

This is one sided test and it is right tailed test

The calculated value of test statisitc is 6.74
degrees of freedom = 7 - 1 = 6

p-value = P(t > 6.74) = 0.0003